1. A 1500 kg car rounds a curve of radius 75 m at a speed of 25 m/s. The curve is banked at an angle of 22° to the horizontal. Calculate:

a) The magnitude of the frictional force

b) The coefficient of static friction



1
Expert's answer
2021-05-14T09:43:06-0400


The forces acting on the wheel:


"F_\text{net}=f\cos\theta+N\sin\theta,\\\space\\
N\cos\theta-mg-f\sin\theta=0,\\\space\\
F_\text{net}=m\frac{v^2}{R}.\\\space\\
m\frac{v^2}{R}=f\cos\theta+N\sin\theta,\\
f=26600\text{ N}.\\
N=32500\text{ N}."

b) The coefficient of friction is


"\mu=\frac fN=0.819."